OpenStudy (mattrushi99):

If f(x)=3x^2-1 g(x)=x+2 find (f+g)(x)

OpenStudy (xguardians):

Take a look at this article and see if it helps: http://www.purplemath.com/modules/fcnops.htm

jhonyy9 (jhonyy9):

do you know what mean (f+g)(x) ?

jhonyy9 (jhonyy9):

@MattRushi99 posted this exercise and is offline

OpenStudy (saracaji98):

isnt it f(x)+g(x)?

jhonyy9 (jhonyy9):

yes is this try solve it for medal - i will give you - ok. ?

jhonyy9 (jhonyy9):

courage this is easy

OpenStudy (saracaji98):

if you add them together you would end up with 3x^2+x+1

jhonyy9 (jhonyy9):

exactly but please write it step by step for the Asker can understanding it easy sure - ok. ?

jhonyy9 (jhonyy9):

in this way f(x)+g(x) = (f+)(x) = ?

jhonyy9 (jhonyy9):

i missed the g from there - sorry

jhonyy9 (jhonyy9):

(f+g)(x)

OpenStudy (saracaji98):

So you have f(x)= 3x^2-1 and g(x)= x+2. since it is asking you to add them, you would have 3x^2-1+x+2. Then, you simplify so you would end up with the answer which is 3x^2+x+1.

jhonyy9 (jhonyy9):

nice job thank you